OpenStudy (amistre64):

a bottle with a cork sells for $1.10. If the bottle cost $1 more than the cork, how much does the bottle cost?

OpenStudy (anonymous):

Is this a trick question? $1 bottle $1.10 bottle w/ cork

OpenStudy (amistre64):

so you propose the cork cost .10; and the bottle costs $1 more than the cork - at $1.10 added together that would be $1.20 right?

OpenStudy (anonymous):

X is cork. Bottle is x + 1.00 . 1.10 = 1.00 + x. Subtract 1.00 and x = .10

OpenStudy (amistre64):

i agree with your x and x+100; but then you dont add those

OpenStudy (amistre64):

lets call it cork = c c + c + 100 = 110 2c + 100 = 110 2c = 110 -100 2c = 10 c = 10/2 c = 5 bottle = 105

OpenStudy (amistre64):

my real question is, why is this in the paradox section of a book i read .... but then i spose this answers it lol

OpenStudy (anonymous):

Why is there 2 c? You are adding two variables of a cork. We are only looking at one bottle, which would have 1 cork.

OpenStudy (amistre64):

the bottle is the price of the cork + $1 right?

OpenStudy (anonymous):

sounds much better when read out loud a bottle and a cork cost a dollar ten cents. i get so high i can't pay the rent

OpenStudy (amistre64):

so its not the number of corks that are bought, but the price that matters

OpenStudy (anonymous):

i love this question. total is $1.10 and the bottle cost $1 more than the cork, so the bottle cost $1.05 and the cork cost $.05

OpenStudy (amistre64):

got a 100 on my journal, but i think its cause i mention credence clearwater revival

OpenStudy (anonymous):

Cork(c) and bottle = 1.10 Bottle is 1.00 more than cork(c) 1.10 - 1.00 = c c+ 1.00 = 1.10 therefore c = .10

OpenStudy (anonymous):

Ohh I see I forget to subtract the c from the total. You're right.

OpenStudy (amistre64):

;)

OpenStudy (anonymous):

i want you to spell new york. a knife a fork a bottle and a cork, that's the way you spell new york http://www.youtube.com/watch?v=8pyB6W8jEWM&feature=related